OpenStudy (anonymous):

Oscar entered y = –0.4^x + 3 into an online graphing tool.

OpenStudy (anonymous):

What function did he graph? (Points : 1) y = (–0.4)x + 3 y = (–0.4)x + 3 y = –(0.4)x + 3 y = –(0.4)x + 3

OpenStudy (anonymous):

this is better